# The dynamics of the Reddit collective action leading to the GameStop short squeeze

### v3

@@ Abstract
In early 2021, the stock prices of GameStop, AMC, Nokia, and BlackBerry experienced dramatic increases, triggered by short squeeze operations that have been largely attributed to Reddit’s retail investors. These events showcased, for the first time, the potential of online social networks to catalyze financial collective action. How, when and to what extent Reddit users played a role in driving up these prices, however, remains unclear. We address these questions by statistical analysis of time series with high temporal resolution, about social activity on Reddit and Twitter as well as stock trading volumes. We find that increasing Reddit discussions anticipated high trading volume before the GameStop short squeeze, with their predictive power being particularly strong on hourly time scales. This effect emerged abruptly a few weeks before the event, but waned once the community of investors gained widespread visibility through Twitter. Meanwhile, the collective investment of the Reddit community, quantified through each user’s financial position on GameStop, closely mirrored the market capitalization of the stock. These evidences suggest a coordinated action by Reddit users in developing a shared financial strategy through social media. Towards the end of January, users talking about GameStop contributed to raise the popularity of BlackBerry, AMC and Nokia, which emerged as the most popular stocks as the community gained global recognition. Overall, our results shed light on the dynamics behind the first large-scale financial collective action driven by social media users.
Introduction
@@
From swinging elections to fueling social movements, instantaneous global communication enabled by online social media can shape leading global phenomena [1, 2, 3, 4, 5]. Alongside sectors like politics, healthcare and education [6, 7, 8], financial markets represent a key area where the activity of social media users can have destabilizing effects with worldwide resonance [9]. Indeed, not only it is by now evident that users activity on mainstream platforms like Twitter [10, 11] and Google Search [12] mirrors very well the state of stock markets. There is also ample evidence that online signals can at times anticipate the evolution of stock prices [10, 13, 12, 14, 15], and thus can be used to assist automated trading [16, 17, 18, 19]. This goes along with the increasingly central role played by non-professional investors in a sector that was once dominated by institutional funds and large corporations [20, 21, 22].
However, the interplay of social media signals and stock market trends is complex and evolving in time. In particular, it mainly emerges around specific events, for example when surges in online interactions catalyze trading activity or, conversely, when large price variations spark online conversation [23]. Thus, while unable to explain long-term market trends, social media signals tend to capture well transient market trends and stock-specific swings [24]. One of the reasons of such connection is that social media are widely used as a medium to exchange financial information [25]. Serving both as the source for real-time decision-making and as a mean for disseminating knowledge and decisions [26, 27, 28], information on social media can at the same time induce, amplify and reflect retail investors’ views of the market [29, 25, 30].
@@
In early 2021, financial markets were shaken by an unprecedented event: the stock of video-game retail company GameStop (GME) experienced a “short squeeze”, with a price surge of almost 1625% within a week [31]. This financial operation was attributed to activity from users of the social media platform Reddit, particularly the subreddit WallStreetBets (WSB), and was rapidly followed by similar market rallies for other stocks: BlackBerry (BB), AMC Entertainment Holdings (AMC) and Nokia (NOK). The GME short squeeze hinted at another way through which social media can impact financial markets: by enabling retail investors to coordinate their investment strategy and challenge the influence of large institutional funds [31]. While this collective action was unprecedented for financial markets, the phenomenon of coordination on social media has been documented in different spheres, from orchestrating on-the-ground protests [32], initiating consumer boycotts [33], to propagating viral challenges for charitable causes [34], and catalyzing social movements [35]. Such “coordination” does not necessarily denote a centralized strategy; instead, it suggests that mere collective discourse might effectively align the actions of individual investors [36, 37, 38, 39]. The Reddit platform could have supported coordination by fostering community-driven discussions [40, 41, 42, 43] focused on shared interests and collective goals. On the contrary, mainstream social media like Twitter are primarily used as news feeds [44], and facilitate coordination only to a limited extent through the use of hashtags [4, 45].
@@
The GME short squeeze has been extensively analyzed. Several studies reported a strong correlation between the activity on WSB and the evolution of the GME market indicators throughout the event [41, 40, 46, 47, 48, 49]. However, two major questions remain unanswered. Firstly, it is still unclear if – and from which specific moment – the activity of WSB users anticipated the actual price surge in GME. Secondly, it is unclear if WSB served merely as a platform for financial discourse or as a hub to coordinate a common investment strategy. In this work, we study the temporal relation and predictability among three key factors in the period around the GME event: the stock market movements, the activity on WSB and the broader public attention to the stocks. Our analysis provides evidence in support of the hypothesis that activity on the WSB community forecast the price surge of GME, BB, AMC and NOK. Further, we shed light on the temporal dynamics of the GME short squeeze, characterizing its evolution through three distinct phases of online behavior: Discussion, Action, and Visibility.
Results and Discussion
Reddit activity on GME mirrors changes of GME trading volume.
@@
We characterise the discourse surrounding the GME event using the volume of GME-related activity on Reddit and on Twitter, the former reflecting the conversation within the community of WSB users and the latter capturing the broader public attention to the stock. Following recent literature [14, 41], we quantify these two signals by measuring the hourly occurrence of the GME ticker in the text of posts and comments on Reddit and Twitter, respectively (see Methods for more details). Starting approximately 15 days before the short squeeze date of January 28th, the GME hourly trading volume and price featured an increasing trend that closely resembles the increase in volume of GME-related conversation on both social networks (see Figure 1a).
@@
To better assess if the financial discourse on Reddit mirrored market indicators, we employ detrended cross-correlation analysis (DCCA) [50]. This method assesses correlations while accounting for prevalent effects observed in financial time series, such as non-linearity, underlying trends, scaling behaviors and long-range dependencies [51]. Specifically, through DCCA, we gauge if fluctuations in GME-related activity are correlated with fluctuations of GME market volume across multiple time scales [14] (see Methods for further details). To capture these long-term dependencies, we consider the absolute value of hourly returns of GME occurrences, stock price and trading volume [50] (see Methods) from December 1st, 2020 to July 1st, 2021, and compute the detrended covariances for each combination of signal pairs (see Figure 1b). First, we observe that such detrended covariances follow power-law scaling (see Methods and Supplementary Materials, Section 3). Remarkably, the GME trading volume displays larger detrended covariance with Reddit activity than with GME price, with a cross-correlation coefficient three times larger than the one observed for the other pairs of signals (see Inset Figure 1b). This suggests that changes in trading volume are more tied to Reddit discussions, rather than to price movements. This is also confirmed by Convergent Cross-Mapping techniques [52] (see Supplementary Materials, Section 5), which demonstrate a synergistic relationship between Reddit activity and trading volume. Convergent Cross Mapping (CCM) quantifies the non-linear coupling between variables within a dynamical system, by assessing whether past states of one variable $x$ can predict future states of another variable $y$, without assuming model-specific dependencies. Overall, we can conclude that large variations in Reddit discussions are strongly related to changes in trading volume, as more discussions correspond to a general increase in buying activity and vice versa.
@@
The collective position of WSB users on GME amounts at  1% of the stock market capitalization.
@@
To capture the actual engagement of WSB users with the GME stock [14, 41], we analyse screenshots of trading activity posted on Reddit [40]. Indeed, some WSB users provide evidence of their financial investments by posting screenshots of their accounts taken from trading platforms, which attest their positions and movements. We analyse such screenshots using computer vision techniques, so to estimate the financial positions of Reddit users on GME [40] (see Methods for more details). We observe that the values of the individual positions on GME follow a Log-normal distribution (see the Inset of Figure 1c). By summing the positions across all users, we obtain the collective investment, which provides a lower-bound estimate of the community stake in GME. We find that this collective position grew significantly 15 days prior to the squeeze. This finding complements the observation that the number of WSB users posting screenshots grew in the period preceding the squeeze [40]. Notably, the growth of the collective position closely follows the market capitalization of GME (see Figure 1c). We estimated that the collective investments of Reddit users represented – on average over the time period considered – at least the 1% of the total market capitalization of GME (see Figure 1c).
Figure 1: The correlation between Reddit activity and GameStop market indicators. a) Daily occurrences of the “GME” ticker in conversations on WSB (orange, y-right axis) and Twitter (lightblue, y-right axis), along with the daily trading volume of the GME stock (blue, right y-axis) and its daily closing price (green, left y-axis). The vertical gray dotted line marks January 13th, identified as the WSB-led action’s start (see subsequent sections), while the dashed line marks January 27th, when Elon Musk’s Tweet broadly publicized the action. To improve the chart’s clarity, we applied a 5-day moving average and normalized the trading volume, Reddit and Twitter signals by their mean. These normalizations were not employed in the analysis. b) Values of detrended variance for price (green), trading volume (blue) and Reddit activity (red) signals, and values of detrended covariance for the pairs price / trading volume (purple), price / Reddit activity (pink), trading volume / Reddit activity (yellow), as a function of the time window considered. Inset: Detrended Cross Correlation Coefficients for paired combinations of GME signals (same color schemes of the main panel). The shaded gray area represents the 95% confidence interval for correlation values obtained from pairs of independent signals. c) Daily value of the WSB collective position on GME (orange, y-right axis) and GME Market Capitalization (blue, y-left axis). Inset: Distribution of values of users’ position on GME (red) alongside the Log-Normal fit (grey line, $\mu=4.96$, $\sigma=3.63$).
@@
Granger analysis identifies three distinct phases of online behavior.
@@
We now test the hypothesis that changes in GME-related activity on social media anticipated changes in GME trading volume around the short squeeze. Specifically, we employ the Granger test to assess the chronological anticipation $x\to y$. The test assesses whether a signal $x$ has predictive power on another signal $y$, building a linear regression model for $y(t)$ leveraging past information on both $y$ and $x$ from $t-\Delta t$ up to $t-1$. The Granger Index [53] quantifies the test result as the ratio of predictive accuracy between the model and a baseline that excludes past information derived from $x$ (see Methods for more details). As $x$ and $y$ we consider the possible pairs of signals among GME trading volume and GME-related discussions on Reddit and Twitter. We consider hourly values during market opening times, within a 15-day time window (corresponding to 120 points per time series – see Methods) between January 5 and February 5. The results are robust to different choices of window length (see Supplementary Materials, Section 4). Before the analysis we test the stationarity of the signals within the specified windows (see Supplementary Materials, Section 4), apply a daily moving average (the results remain robust when we do not apply the moving average, see Supplementary Materials, Section 4) and compute the hourly log returns.
Our analysis of the Granger Index for $\Delta t=1$, meaning that the model uses $x(t-1)$ to predict $y(t)$, highlights three distinct phases of online behavior (see Figure 2a).
@@
Before January 13th, the Granger Index is not significantly different from zero (p-values $>0.1$) for all combinations of signals, meaning that no anticipatory effects are observed.
@@
On January 13, the Granger Index capturing the anticipatory power of the Reddit activity on trading volume has a sharp transition to a very high value of 0.47 (p-values $<0.1$). Simultaneously, also the predictive power of the trading volume on Reddit becomes significant; however, the Granger Index in the Reddit-to-Trading Volume direction is roughly three times larger than in the opposite Trading Volume-to-Reddit direction. These Indices remain significant – although steadily declining – for about 12 days.
@@
On January 27th, the Tweet “Gamestonk!!” by Elon Musk brought the GME short squeeze to the public attention. On this date, the anticipatory power of Reddit activity on the trading volume (and viceversa) becomes non significant. Instead, Reddit activity starts to anticipate the activity on Twitter: once they became informed about the short squeeze, Twitter users turned to Reddit for information about GME.
@@
The analysis in the previous section considers only relations between pairs of signals, neglecting the role played by the interactions between the three of them. Further, the identified anticipatory power of Reddit activity on the Trading Volume may have either a reinforcing nature, where for instance an increase in social media activity anticipates an increase of market activity, or alternatively, a counteracting effect. To discern these scenarios, as well as to capture how the relationship between the three signals under study change over time, here we employ a Multivariate Vector Autoregressive model. The method entails fitting a linear model that predicts a target output signal $y$ at time $t$, using past information from $t-1$ to $t-\Delta t$ on $y$ itself plus one or more input feature signals $\{x_{1},x_{2},x_{3},\dots\}$. The coefficients of the model capture how the individual input signals contribute to predicting future values of $y$ (see Methods for more details). In particular, a positive coefficient for $x_{i}$ indicates a positive reinforcing relationship of $x_{i}$ on $y$, while a negative coefficient an opposite and counteracting effect. We apply the method to study the hourly log-returns of Twitter activity, Reddit activity, and trading volume. As we did for the Granger analysys, we fitted the model within each 15-day period ending on a day between January 5th and February 5th, for $\Delta t=1$ (results for other window period lengths are reported in the Supplementary Materials, Section 4).
@@
Figure 2b shows the coefficients that capture the predictive power of the three signals on the future values of Twitter (top plot), Reddit (mid plot), and trading volume (bottom plot). First, we observe a self-reinforcing dynamic within Reddit and Twitter throughout the whole period under study, where increased activity within the social network leads to further increases. This is evident from the consistently positive and significant coefficients of the Reddit-to-Reddit and Twitter-to-Twitter relation. During the Action phase, the self-reinforcing nature of Reddit is even more pronounced.
@@
Secondly, during the Action phase, the coefficient capturing the Reddit-to-Trading Volume relation becomes significant and positive, implying that the increase of activity on Reddit is followed – within the next hour – by an increase in trading volume. Notably, in this phase the magnitude of the Reddit-to-Trading Volume coefficient is similar to the magnitude of the Reddit-to-Reddit coefficient (the two can be directly compared since we are considering the log-returns of the signals).
@@
Finally, during the Visibility phase, the trading volume starts following a self-reinforcing dynamics (as indicated by the significant and positive coefficient for the Trading Volume-to-Trading Volume relation), while the Reddit-to-Trading Volume coefficient becomes non significant. In this phase, the Reddit-to-Twitter coefficient aligns in both magnitude and direction with the Twitter-to-Twitter coefficient, again highlighting that public attention was directed towards WSB.
@@
By conducting this analysis at various lags $\Delta t$, we found that the coefficients at $\Delta t=1$ hour are larger than the others (see Supplementary Materials, Section 4), indicating the greater significance of short time scales in the entwined dynamics of social media and markets. Interestingly, the typical reply speed, defined as the average time interval between a comment and its direct responses, varies between $\sim 2.5$ hours during the Discussion phase to $\sim 0.5$ hours during the Visibility phase (see Figure 2c). Hence, the influence of Reddit on the market occurs on a timescale closely aligned with that of the unfolding conversations.
Figure 2: Reddit activity anticipates trading volume during the short squeeze. In all the following subplots the vertical gray dotted line indicates the beginning of the Action phase (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021) that brought the squeeze to the public attention. Triangles correspond to statistical significant values (p-values $<0.1$). a) Granger Index capturing the predictive power of a signal on another (with a lag of 1 hour) for the following pairs: Trading Volume-to-Reddit (blue), Reddit-to-Trading Volume (red), Trading Volume-to-Twitter (gray) and Reddit-to-Twitter (yellow). Each point is computed considering time series spanning the 15 preceding days. b) Each panel shows the coefficients of a Multivariate Vector Autoregressive Model predicting Twitter (top), Reddit (middle), and Trading Volume (bottom) activities, corresponding to antecedent values of Reddit (solid orange), Twitter (dashed lightblue), and trading volume (dotted blue). Each point is computed considering time series spanning the 15 preceding days. c) Daily average (and standard deviation of the mean) of the reply speed, defined as the average time interval between a a comment and its direct responses under a Reddit post. The average by post is weighted using the number of comments underneath the post. We apply a 5-day moving average to the time series and consider only weekdays to enhance the chart’s clarity.
@@
The Granger analysis reveals that, during the GME Action phase, the Reddit-to-Trading Volume anticipatory relation is only significant for BB (see Figure 3b). Interestingly, the Granger Index for BB becomes statistically significant only five days after it does so for GME. The analysis for AMC and NOK, instead, does not lead to significant conclusions (see Supplementary Materials, Section 4), possibly due to the less attention they received on Reddit (see Figure 3a). We refer the reader to the Supplementary Materials, Section 3 for the analyses of long-range detrended cross-correlations (DCCA) and Convergent Cross-Mapping between trading volume and the frequency of occurrences for BB, AMC and NOK.
@@
Figure 3: BB, AMC, NOK and the role of GME-engaged users. In all the following subplots the vertical gray dotted line indicates the beginning of the GME Action (13 January 2021), whereas the vertical gray dashed line corresponds to the Tweet by Elon Musk (27 January 2021). Triangles refer to statistically significant values (p-value < 0.1). a) (top) Rank chart of stock tickers on WSB based on their average daily occurrence within each 15-days window. a) (bottom) Cumulative number of occurrences of a given stock on WSB. b) Granger Index for possible combinations of BB, NOK, AMC signals. c) Coefficients of the Multivariate Vector Autoregressive Model for BB-related activity on Reddit (top panel) and BB trading volume (bottom panel), related to the Reddit (solid orange) and trading volume (dashed blue) signals. d) Overlap index for pairs of stocks, computed over a 5-day window with a one-day shift, measuring the proportion of users discussing each pair of stocks. We observe in both panels an increasing overlap after the starting of the GME rally. e) Jaccard index for pairs of stocks, computed over a 5-day window with a one-day shift, measuring the proportion of users discussing both pair of stocks.
Conclusion

### v4

@@ Abstract
In early 2021, the stock prices of GameStop, AMC, Nokia and BlackBerry experienced dramatic increases, triggered by short squeeze operations that have been largely attributed to Reddit’s retail investors. Here we shed light on the extent and timing of Reddit users’ influence on the GameStop short squeeze. Using statistical analysis tools with high temporal resolution, we find that increasing Reddit discussions anticipated high trading volumes. This effect emerged abruptly a few weeks before the event but waned once the community gained widespread visibility through Twitter. Meanwhile, the collective investment of the community, quantified through posts of individual positions, closely mirrored the market capitalization of the stock. This evidence suggests a coordinated action of users in developing a shared financial strategy through social media – targeting GameStop first and other stocks afterwards. Overall, our results provide novel insights into the role of Reddit users in the dynamics of the GameStop short squeeze.
1 Introduction
@@
From swinging elections to fueling social movements, instantaneous global communication enabled by online social media can shape leading global phenomena [1, 2, 3, 4, 5]. Financial markets represent a key area where the activity of social media users can have destabilizing effects with worldwide resonance [6]. Indeed, there is ample evidence that user activity on mainstream platforms like Twitter [7, 8] and Google Search [9] not only mirror the mood of the market but can at times anticipate the evolution of stock prices [7, 10, 9, 11, 12, 13]. This goes along with the increasingly central role played by retail investors, or non-professional investors, in a sector that was once dominated by institutional funds and large corporations [14, 15, 16].
The dissemination of knowledge on social media stems directly from the dynamic interactions of users [17], who actively share knowledge and opinions on financial markets[18]. Thus, this medium has the potential to enhance the discovery of information that underlies price formation in financial markets [19]. Social communication, indeed, supports the price-discovery process [20] and improves market efficiency, when information is exogenous [21]. However, more than rational agents, social media users tend to resemble noisy traders [22], suggesting that standard models of efficient markets should account for behavioral factors, such as investors’ sentiment, public mood [23], cultural traits, psychological biases, social network structure, and information asymmetries [24] as influential for price formation. What emerges from this picture is the complex and evolving interplay between social media signals and stock market trends, which enables narratives, ideas, and investment strategies to rapidly spread to a wide audience [25, 26]. Notably, this relationship mainly emerges around specific events, for example when surges in online interactions catalyze trading activity or, conversely, when large price variations spark online conversation [27]. As such, social media signals tend to capture well transient market trends and stock-specific swings [28].
@@
In early 2021, financial markets were shaken by an unprecedented event: the stock of video-game retail company GameStop (GME) experienced a “short squeeze”, with a price surge of almost 1625% within a week [29]. This financial operation was attributed to activity from users of the social media platform Reddit, particularly the subreddit WallStreetBets (WSB), and was rapidly followed by similar market rallies for other stocks: BlackBerry (BB), AMC Entertainment Holdings (AMC) and Nokia (NOK). The GME short squeeze suggested that retail investors, by sharing their investment strategies on social media, could potentially challenge the influence of large institutional funds [29]. The Reddit platform could have supported coordination by fostering community-driven discussions [30, 26, 31, 32] focused on shared interests and collective goals. On the contrary, mainstream social media like Twitter are primarily used as news feeds [33], and facilitate coordination only to a limited extent through the use of hashtags [4, 34].
@@
The GME short squeeze has been extensively analyzed in the scientific literature. Several studies reported a strong correlation between the activity on WSB and the evolution of the GME price throughout the event [26, 30, 35, 36, 37, 38]. However, two major questions remain unanswered regarding the role of WSB users. Firstly, it is still unclear if – and from which specific moment – the activity of WSB users anticipated the actual GME price surge. Secondly, it is unclear if WSB served merely as a platform for financial discourse or as a hub to coordinate a common investment strategy. In this work, we study the temporal relation among three key signals in the period around the GME event: the stock market movements, the activity on WSB and the broader public attention to the stocks. We employ different techniques, from standard Granger Analysis and Multivariate Vector Autoregressive Models to study linear effects, to Detrended Cross-Correlation Analysis (DCCA) and Convergent Cross-Mapping (CCM) to capture the non-linear and complex character of the phenomenon. Our analysis provides evidence in support of the hypothesis that activity on the WSB community forecast the price surge of GME, BB, AMC and NOK. Further, we shed light on the temporal dynamics of the GME short squeeze, characterizing its evolution through three distinct phases of online behavior: Discussion, Action, and Visibility.
2 Results
2.1 Reddit activity on GME mirrors changes of GME trading volume.
@@ 2 Results
We characterise the discourse during the GameStop saga using the volume of GME-related activity on Reddit and on Twitter (see Supplementary Materials, Section 1), the former reflecting the conversation within the community of WSB users and the latter capturing the broader public attention to the stock. Following recent literature [11, 26], we quantify these two signals by measuring the hourly occurrence of the GME ticker in the text of posts and comments on Reddit and Twitter, respectively (see Methods for more details). The GameStop saga is described by the series of events reported in Table 1. Interest in GME within the WSB community began to surge as early as December 2020 [30, 26]. However, despite this growing public discourse, the hourly trading volume and price of GME did not exhibit a significant upward trend until approximately 15 days before the short squeeze date of January 27th – which we associated with Elon Musk’s tweet. As shown in Figure 1a, this trend closely resembles the increase in GME-related conversation volume on social media.
@@
This observation suggests a deeper connection between the two signals. To shed light on this point, we employ detrended cross-correlation analysis (DCCA). This method allows investigating cross-correlations between time series in the presence of non-stationarity [39], hence addressing prevalent effects observed in financial time series such as non-linearity, long-term underlying trends and long-range dependencies [40]. Specifically, through DCCA, we gauge if fluctuations in GME-related activity are correlated with fluctuations of GME market volume across multiple time scales [11] (see Methods for further details). We apply DCCA to time series of absolute values of hourly returns for GME occurrences, stock price and trading volume [39] from December 1st, 2020 to July 1st, 2021, computing the detrended covariances for each combination of signal pairs (see Methods). Results shown in Figure 1b indicate that such detrended covariances follow power-law scaling, pointing to long-term coupling between the signals (see Supplementary Materials, Section 2). Remarkably, the GME trading volume displays larger detrended covariance with Reddit activity than with GME price, with a cross-correlation coefficient three times larger than the one observed for the other pairs of signals (see Inset Figure 1b). This suggests that changes in trading volume are more tied to Reddit discussions, rather than to price movements.
We further employ Convergent Cross-Mapping (CCM) to corroborate the dynamic interdependence between Reddit activity and trading volume [41]. This method quantifies the non-linear coupling between variables by assessing whether past states of one variable $x$ can predict future states of another variable $y$, without assuming model-specific dependencies. Results of this approach confirm a strong dynamic coupling between Reddit activity and trading volume (see Supplementary Materials, Section 3). Overall, we can conclude that large variations in Reddit discussions are strongly related to changes in trading volume, as more discussions correspond to a general increase in buying activity and vice versa.
@@
2.2 The collective position of WSB users on GME amounts to at least 1% of the stock’s market capitalization.
@@
In WSB, some users provide evidence of their financial investments by posting screenshots of their accounts taken from trading platforms, which attest their positions and trading activity [30]. We analyse such screenshots to capture the actual engagement of WSB users with the GME stock [11, 26], using computer vision techniques (see Methods for more details and Supplementary Materials, Section 4). We observe that the values of the individual positions on GME follow a Log-normal distribution (see the Inset of Figure 1c). By summing the positions across all users, we obtain the posted collective investment, which provides a lower-bound estimate of the community stake in GME, as it includes only the retail investors who have publicly shared their positions. We find that this posted collective position grew significantly 15 days prior to the squeeze. This finding complements the observation that the number of WSB users posting screenshots grew in the period preceding the squeeze [30], likely indicating that we have more users investing in GME. Notably, the growth of the posted collective position closely follows the market capitalization of GME (see Figure 1c). We estimate that the posted collective investments of Reddit users represented – on average over the time period considered – at least the 1% of the total market capitalization of GME (see Figure 1c).
@@
2.3 Granger analysis identifies three distinct phases of online behavior.
@@
We now test the hypothesis that changes in GME-related activity on social media anticipated changes in GME trading volume around the short squeeze. Specifically, we employ the Granger test to assess the chronological anticipation $x\to y$: whether a signal $x$ has predictive power on another signal $y$, building a linear regression model for $y(t)$ leveraging past information on both $y$ and $x$ from $t-\Delta t$ up to $t-1$. The Granger Index [42] quantifies the test result as the ratio of predictive accuracy between the model and a baseline that excludes past information derived from $x$ (see Methods for more details). As $x$ and $y$ we consider the possible pairs of signals among GME trading volume and GME-related discussions on Reddit and Twitter. We consider hourly values during market opening times, within a 15-day time window (corresponding to 120 points per time series – see Methods) between January 5 and February 5. This choice of window size ensures that we capture short-term patterns that are relevant to our case study [28] and at the same time that we have enough statistics to draw robust conclusions from the Granger analyses, even in regimes of small coupling [43]. Results are, however, robust to different choices of window length (see Supplementary Materials, Section 5). After testing the stationarity of the time series within the specified windows (see Supplementary Materials, Section 5), we apply a daily moving average (the results remain robust when we do not apply the moving average, see Supplementary Materials, Section 5) and compute the hourly log returns of the signals.
Figure 2a shows the resulting Granger Index for $\Delta t=1$, meaning that the model uses $x(t-1)$ to predict $y(t)$. In the plot, triangles correspond to statistical significant values (p-values $<0.05$). Size effects are instead captured by the coefficient of the regression of the independent variable, reported in Figure 2b for the Reddit-to-Trading Volume direction (see Supplementary Materials, Section 5 for other values). Results highlight three distinct phases of online behavior (see Figure 2a).
@@
Before January 13th, the Granger Index and size effects are not significantly different from zero (p-values $>0.1$) for all combinations of signals, meaning that no anticipatory effects are observed.
@@
On January 13, the Granger Index capturing the anticipatory power of the Reddit activity on trading volume has a sharp transition to a very high value of 0.47 (p-values $<0.05$). At the same time, size effects grow up to 1.5, meaning that social media activity has a large positive effect in the prediction of market trends. Simultaneously, the predictive power of the trading volume on Reddit is significant; however, the Granger Index in the Reddit-to-Trading Volume direction is roughly three times larger than in the opposite Trading Volume-to-Reddit direction. These indices remain significant – although steadily declining – for about 12 days. The p-values are reported in Supplementary Materials, Section 5.
@@
On January 27th, the Tweet “Gamestonk!!” by Elon Musk brought the GME short squeeze to the public attention. On this date, the anticipatory power of Reddit activity on the trading volume (and viceversa) is non significant, and size effects vanish. Instead, Reddit activity starts to anticipate the activity on Twitter: once they became informed about the short squeeze, Twitter users turned to Reddit for information about GME.
@@
The analysis in the previous section considers only pairwise relations among signals. In order to capture how the mutual relationship between all three signals under study change over time, we employ a Multivariate Vector Autoregressive model. The method generalises the Granger analysis to multiple signals: it entails fitting a linear model that predicts a target output signal $y$ at time $t$, using past information from $t-1$ to $t-\Delta t$ on $y$ itself plus one or more input feature signals $\{x_{1},x_{2},x_{3},\dots\}$. The coefficients of the model capture how the individual input signals contribute to predicting future values of $y$ (see Methods for more details). In particular, a positive coefficient for $x_{i}$ indicates a positive reinforcing relationship of $x_{i}$ on $y$, while a negative coefficient an opposite and counteracting effect. We apply the method to study the hourly log-returns of Twitter activity, Reddit activity, and trading volume. As we did for the Granger analysys, we fitted the model within each 15-day period ending on a day between January 5th and February 5th, for $\Delta t=1$ (results for other window period lengths are reported in the Supplementary Materials, Section 5).
@@
Figure 2c shows the coefficients that capture the predictive power of the three signals on the future values of Twitter (top plot), Reddit (mid plot), and trading volume (bottom plot). First, we observe a self-reinforcing dynamic within Reddit and Twitter throughout the whole period under study, where increased activity within the social network leads to further increases. This is evident from the consistently positive and significant coefficients of the Reddit-to-Reddit and Twitter-to-Twitter relation. During the Action phase, the self-reinforcing nature of Reddit is even more pronounced.
@@
Secondly, during the Action phase, the coefficient capturing the Reddit-to-Trading Volume relation is significant and positive, implying that the increase of activity on Reddit is followed – within the next hour – by an increase in trading volume. This is coherent with the size effects observed in the Granger analysis. Notably, in this phase the magnitude of the Reddit-to-Trading Volume coefficient is similar to the magnitude of the Reddit-to-Reddit coefficient (the two can be directly compared since we are considering the log-returns of the signals).
@@
Finally, during the Visibility phase, the trading volume starts following a self-reinforcing dynamics (as indicated by the significant and positive coefficient for the Trading Volume-to-Trading Volume relation), while the Reddit-to-Trading Volume coefficient is non significant. In this phase, the Reddit-to-Twitter coefficient aligns in both magnitude and direction with the Twitter-to-Twitter coefficient, again highlighting that public attention was directed towards WSB.
@@
By conducting this analysis at various lags $\Delta t$, we found that the coefficients at $\Delta t=1$ hour are larger than the others (see Supplementary Materials, Section 5), indicating the greater significance of short time scales in the entwined dynamics of social media and markets. Interestingly, the typical reply speed, defined as the average time interval between a comment and its direct responses, varies between $\sim 2.5$ hours during the Discussion phase to $\sim 0.5$ hours during the Visibility phase (see Supplementary Materials, Section 6). Hence, the influence of Reddit on the market occurs on a timescale closely aligned with that of the unfolding conversations.
@@
The Granger analysis reveals that, during the GME Action phase, the Reddit-to-Trading Volume anticipatory relation is only significant for BB (see Figure 3b). Interestingly, the Granger Index for BB is statistically significant only five days after it does so for GME. The analysis for AMC and NOK, instead, does not lead to significant conclusions (see Supplementary Materials, Section 5), possibly due to the less attention they received on Reddit (see Figure 3a). We refer the reader to the Supplementary Materials, Section 2 for the analyses of long-range detrended cross-correlations (DCCA) and Convergent Cross-Mapping (Supplementary Materials, Section 3) between trading volume and the frequency of occurrences for BB, AMC and NOK.
@@
3 Discussion